# Efficient Prior Selection in Gaussian Process Bandits with Thompson Sampling

### v1

@@ A.2 HP-GP-TS
Combining the bounds for (1) and (2), we obtain
${\text{BR}}(T)\leq\frac{\pi^{2}}{6}+\sqrt{CT\beta_{t}{\mathbb{E}}_{p^{*}}[\gamma_{T,p^{*}}]}.$ (60)
Appendix B Additional experimental details
In this section, we provide some additional experimental details and results. First, we include the full set of confusion matrices for the synthetic lengthscale and hill experiments in Fig. 10. In the lengthscale experiments, we observe that PE-GP-UCB and -TS oversample the shortest lengthscale. This is similar to the kernel experiment where the Matérn 3/2 kernel was also oversampled. However, we see that HP-GP-TS and MAP GP-TS do not suffer from this optimistic bias. In the hill experiment, we see that PE-GP-UCB almost never selects Hill 0, i.e. the mean function with only smaller hills. PE-GP-TS selects the prior almost uniformly whilst HP-GP-TS and MAP-GP-TS select the correct mean function in a majority of the time steps.
In Fig. 11, we visualize the median cumulative regret on the real-world data experiments.
@@ Appendix B Additional experimental details
In the Intel experiment, we removed one outlier seed. All methods had a final cumulative regret around 6000°C, note that the average for the worst performing model across the other seeds was $\approx 150$°C. The outlier is shown Fig. 12. We can see that one of the sensors display very high temperatures compared to all other sensors, which is why all methods performed poorly on this seed. It should be noted that many of the sensors in the Intel data logged degrees above 100°C after a certain time - likely due to sensor failure rather boiling temperatures in an office environment. Also note that these days were excluded from both our training and test data. The outlier could be indication that this particular sensor was starting to fail earlier than others.
(a) Lengthscale experiment
(b) Hill experiment
Figure 10: Confusion matrices for the true prior $p^{*}$ and $p_{t}$ across all time steps of the synthetic lengthscale and hill experiments.
Figure 11: Median cumulative regret on Intel temperature data (top) and PeMS data (bottom). Errorbars correspond to first and last decile.
@@
Figure 12: Removed sample from the test data in the Intel experiment. One of the sensors display very high temperatures but

### v2

@@ A.4 Information-theoretic regret bound for GP-TS
Let $Z=f(x)+\epsilon_{t}-{\mathbb{E}}_{t}[f(x)]$. Note that $Z|H_{t}$ is $\sqrt{\sigma_{0}^{2}+\sigma^{2}}$-subgaussian. Consequently, $\log{\mathbb{E}}_{t}[\exp(\lambda Z)]\leq\frac{\lambda^{2}(\sigma_{0}^{2}+\sigma^{2})}{2}$ $\forall\lambda\in{\mathbb{R}}$.
Fix $x^{*},x\in{\mathcal{X}}$. Then,
$\lambda\Big({\mathbb{E}}_{t}[f(x)+\epsilon_{t}|x^{*}=x^{*}]-{\mathbb{E}}_{t}[f(x)]\Big)-\frac{\lambda^{2}(\sigma_{0}^{2}+\sigma^{2})}{2}$ (90)
$\leq\lambda\Big({\mathbb{E}}_{t}[f(x)+\epsilon_{t}|x^{*}=x^{*}]-{\mathbb{E}}_{t}[f(x)]\Big)$ (91)
$\quad-\log{\mathbb{E}}_{t}[\exp(\lambda(f(x)+\epsilon_{t}-{\mathbb{E}}_{t}[f(x)])]$ $(Z|H_{t}\text{ is }\sqrt{\sigma_{0}^{2}+\sigma^{2}}\text{-subgaussian})$ (92)
$=\lambda{\mathbb{E}}_{t}[Z|x^{*}=x^{*}]-\log{\mathbb{E}}_{t}[\exp(\lambda Z)]$ (93)
@@
Now, let $\lambda=\frac{1}{\sigma_{0}^{2}+\sigma^{2}}\left({\mathbb{E}}_{t}[f(x)+\epsilon_{t}|x^{*}=x^{*}]-{\mathbb{E}}_{t}[f(x)]\right)$, then the following holds for all $x^{*},x\in{\mathcal{X}}$
${\mathbb{E}}_{t}\left[f(x)|x^{*}=x^{*}\right]-{\mathbb{E}}_{t}[f(x)]\leq\sqrt{2(\sigma_{0}^{2}+\sigma^{2})D\left({\mathbb{P}}_{t}(f(x)+\epsilon_{t}|x^{*}=x^{*})||{\mathbb{P}}_{t}(f(x)+\epsilon_{t})\right)}.$ (95)
Let $I_{t}(\cdot\,;\cdot\,)=I(\cdot\,;\cdot\,|H_{t}=h_{t})$. Then,
$I_{t}(x^{*};(x_{t},y_{t}))=I_{t}(x^{*};x_{t})+I_{t}(x^{*};y_{t}|x_{t})$ $(\text{Chain rule})$ (96)
$=I_{t}(x^{*};y_{t}|x_{t})$ $(x^{*}|H_{t}{\perp\!\!\!\!\perp}x_{t}|H_{t})$ (97)
$=\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x_{t}=x)I_{t}(x^{*};y_{t}|x_{t}=x)$ (98)
$=\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x_{t}=x)I_{t}(x^{*};f(x_{t})+\epsilon_{t}|x_{t}=x)$ (99)
$=\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x_{t}=x)I_{t}(x^{*};f(x)+\epsilon_{t})$ $(f,x^{*}|H_{t}{\perp\!\!\!\!\perp}x_{t}|H_{t})$ (100)
$=\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x_{t}=x)\Bigg(\sum_{x^{*}\in{\mathcal{X}}}{\mathbb{P}}_{t}(x^{*}=x^{*})\cdot$ (101)
$\quad\quad D\left({\mathbb{P}}_{t}(f(x)+\epsilon_{t}|x^{*}=x^{*})\,\Big|\Big|\,{\mathbb{P}}_{t}(f(x)+\epsilon_{t})\right)\Bigg)$ (102)
$=\sum_{x,x^{*}\in{\mathcal{X}}}{\mathbb{P}}_{t}(x^{*}=x){\mathbb{P}}_{t}(x^{*}=x^{*})\cdot$ $(x_{t}|H_{t}\overset{d}{=}x^{*}|H_{t})$ (103)
$\quad\quad D\left({\mathbb{P}}_{t}(f(x)+\epsilon_{t}|x^{*}=x^{*})\,\Big|\Big|\,{\mathbb{P}}_{t}(f(x)+\epsilon_{t})\right)$ (104)
Putting the above together, we now bound ${\mathbb{E}}_{t}[f(x^{*})-f(x_{t})]^{2}$:
${\mathbb{E}}_{t}[f(x^{*})-f(x_{t})]^{2}=\left(\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x^{*}=x)\left({\mathbb{E}}_{t}[f(x)|x^{*}=x]-{\mathbb{E}}_{t}[f(x)]\right)\right)^{2}$ (105)
$\leq|{\mathcal{X}}|\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x^{*}=x)^{2}\left({\mathbb{E}}_{t}[f(x)|x^{*}=x]-{\mathbb{E}}_{t}[f(x)]\right)^{2}$ $(\text{Cauchy-Schwarz})$ (106)
$\leq|{\mathcal{X}}|\sum_{x,x^{*}\in{\mathcal{X}}}{\mathbb{P}}_{t}(x^{*}=x){\mathbb{P}}_{t}(x^{*}=x^{*})\left({\mathbb{E}}_{t}[f(x)|x^{*}=x^{*}]-{\mathbb{E}}_{t}[f(x)]\right)^{2}$ (107)
$\leq 2|{\mathcal{X}}|(\sigma^{2}_{0}+\sigma^{2})\sum_{x,x^{*}\in{\mathcal{X}}}\Big({\mathbb{P}}_{t}(x^{*}=x){\mathbb{P}}_{t}(x^{*}=x^{*})\cdot$ (108)
$\qquad\qquad\qquad\qquad\qquad D\left({\mathbb{P}}_{t}(f(x)+\epsilon_{t}|x^{*}=x^{*})||{\mathbb{P}}_{t}(f(x)+\epsilon_{t})\right)\Big)$ (109)
$\leq 2|{\mathcal{X}}|(\sigma^{2}_{0}+\sigma^{2})I_{t}(x^{*};(x_{t},y_{t})).$ (110)
Returning to the full regret,
${\mathbb{E}}\bigg[{\sum_{t\in[T]}}f(x^{*})-f(x_{t})\bigg]$ $={\mathbb{E}}\left[{\sum_{t\in[T]}}\sqrt{{\mathbb{E}}_{t}\left[f(x^{*})-f(x_{t})\right]^{2}\cdot\frac{I_{t}(x^{*}\,;(x_{t},y_{t}))}{I_{t}(x^{*}\,;(x_{t},y_{t}))}}\right]$ (111)
$\leq\sqrt{2|{\mathcal{X}}|(\sigma^{2}_{0}+\sigma^{2})}{\mathbb{E}}\left[{\sum_{t\in[T]}}\sqrt{I_{t}(x^{*}\,;(x_{t},y_{t}))}\right]$ (112)
Then, the expectation can then be bounded as
${\mathbb{E}}\bigg[{\sum_{t\in[T]}}\sqrt{I_{t}(x^{*}\,;(x_{t},y_{t}))}\bigg]=$ $\leq{\sum_{t\in[T]}}{\mathbb{E}}\left[\sqrt{I_{t}(x^{*}\,;(x_{t},y_{t}))}\right]$ (114)
$\leq{\sum_{t\in[T]}}1\cdot\sqrt{{\mathbb{E}}\left[I_{t}(x^{*}\,;(x_{t},y_{t}))\right]}$ $\left(\text{Jensen's inequality}\right)$ (115)
$={\sum_{t\in[T]}}\sqrt{{\mathbb{E}}_{H_{t}}\left[I(x^{*}\,;(x_{t},y_{t})|H_{t}=h_{t})\right]}$ (116)
$={\sum_{t\in[T]}}\sqrt{I(x^{*}\,;(x_{t},y_{t})|H_{t})}$ (117)
$\leq\sqrt{{\sum_{t\in[T]}}1^{2}{\sum_{t\in[T]}}I(x^{*}\,;(x_{t},y_{t})|H_{t})}$ $(\text{Cauchy-Schwarz})$ (118)
Finally, the regret can be bounded as
${\mathbb{E}}\bigg[{\sum_{t\in[T]}}f(x^{*})-f(x_{t})\bigg]$ $=\sqrt{2|{\mathcal{X}}|(\sigma^{2}_{0}+\sigma^{2})}\sqrt{TI(x^{*};H_{T})}$ (119)
$\leq\sqrt{2|{\mathcal{X}}|(\sigma^{2}_{0}+\sigma^{2})}\sqrt{TH(x^{*})}$ (120)
$\leq\sqrt{2|{\mathcal{X}}|(\sigma^{2}_{0}+\sigma^{2})}\sqrt{T\log|{\mathcal{X}}|}$ (121)
Appendix B Description of kernels
[…]
Appendix C Additional experimental details
In this section, we provide some additional details about the experiments.
For all the real-world datasets, sensors containing any null measurements have been filtered out.
The Intel Berkeley dataset consists of measurements from 46 temperature sensors across 19 days. The training set consists of the first 12 days of measurements and the remaining 7 days constitute the test set. The noise variance is set to $\sigma^{2}=0.7^{2}$.
The PeMS data consists of measurements from 211 sensors along the I-880 highway from all of 2023. The goal is to find the sensors with low speeds to identify congestions. We use the 5-min averages provided by PeMS. Data between 2023-01-01 and 2023-09-01 is put into the training set whilst the data until 2023-12-31 is put into the test set. The noise variance is set to $\sigma^{2}=2.25^{2}$.
The PNW precipication data consists of daily precipication data from 1949 to 1950 across 167 $50\times 50$ km regions in the Pacific Northwest. The goal is to find the region with the highest precipitation for any given day. The training data consists of the measurements made prior to 1980 and the test data consists of the measurements between 1980 and 1994. The original data is stated to be given in mm/day however the data seems to be off by a factor of 10. We rescale the data to a log-scale using $\log(\cdot/10+0.1)$, similar to Krause et al. (2008). The noise variance is set to $\sigma^{2}=0.41^{2}$.
@@ Appendix C Additional experimental details
In the Intel experiment, we removed one outlier seed. All methods had a final cumulative regret around 6000°C, note that the average for the worst performing model across the other seeds was $\approx 150$°C. The outlier is shown Fig.˜8. We can see that one of the sensors display very high temperatures compared to all other sensors, which is why all methods performed poorly on this seed. It should be noted that many of the sensors in the Intel data logged degrees above 100°C after a certain time - likely due to sensor failure rather boiling temperatures in an office environment. Also note that these days were excluded from both our training and test data. The outlier could be an indication that this particular sensor was starting to fail earlier than others.
@@
Figure 8: Removed sample from the test data in the Intel experiment. One of the sensors display very high temperatures.
Appendix D Additional experimental results
In this section, we provide some additional experimental results.
Figure 9: Mean number of priors remaining in $P_{t}$ over time for PE-GP-UCB and -TS.
Figure 10: Average entropy in the hyperposterior $P_{t}$ over time for HP- and MAP GP-TS. The dashed reference values correspond to entropies of discrete distributions with prob. $q$ on one choice and prob. $\frac{1-q}{|P|-1}$ on the other |P|1|P|-1 choices.
First, we include the mean number of priors in $P_{t}$ for all experiments in Fig.˜9. Similarly, we include the average entropy of the hyperposterior for all experiments in Fig.˜10. For the lengthscale, subspace, PeMS and PNW precipication experiments, hardly any priors are eliminated. In contrast, the hyperposterior entropy concentrates rapidly across all experiments with the subspace and PNW precipication having the most and least concentrated hyperposterior.
We include the full set of confusion matrices for the lengthscale and subspace experiments in Fig.˜11. In the lengthscale experiments, we observe that PE-GP-UCB and -TS oversample the shortest lengthscale. This is similar to the kernel experiment where the Matérn 3/2 kernel was also oversampled. However, we see that HP-GP-TS and MAP GP-TS do not suffer from this optimistic bias. In the subspace experiment, HP- and MAP GP-TS have an accuracy of around 96% where as PE-GP-TS and -UCB have accuracies 30% and 36% respectively. The priors are equivalent up to coordinate permutations and are therefore difficult. The PE-methods do not oversample any specific prior but commit to much time to exploring along the irrelevant dimensions.
In Tables˜1 and 2, the total regret for the lengthscale and subspace scaling experiments are shown.
In Fig.˜12, we visualize the median cumulative regret on the real-world data experiments. Similarly, in Fig.˜13, we show further quantiles of the final cumulative regret. MAP and HP-GP-TS exhibit similar median regret across the experiments but MAP GP-TS has a larger variance and tail values.
(a) Kernel experiment
(b) Lengthscale experiment
(c) Subspace experiment
Figure 11: Confusion matrices for the true prior $p^{*}$ and $p_{t}$ across all time steps of the synthetic experiments.
Table 1: Average total regret and $\pm 1$ standard error for the lengthscale experiment as $|P|$ increases.
Algorithm Lengthscales, $|P|$
8 16 32 64 128
MAP GP-TS 30.2 $\pm$ 1.2 32.4 $\pm$ 2.5 32.5 $\pm$ 2.1 28.7 $\pm$ 1.1 30.8 $\pm$ 1.9
HP-GP-TS 31.4 $\pm$ 1.0 31.7 $\pm$ 0.9 30.8 $\pm$ 0.8 30.7 $\pm$ 1.0 31.0 $\pm$ 1.4
PE-GP-TS 61.8 $\pm$ 0.5 61.3 $\pm$ 0.5 62.2 $\pm$ 0.5 62.4 $\pm$ 0.4 64.3 $\pm$ 0.4
PE-GP-UCB 114.2 $\pm$ 0.6 114.8 $\pm$ 0.6 115.5 $\pm$ 0.6 114.5 $\pm$ 0.6 114.8 $\pm$ 0.6
Oracle GP-TS 28.1 $\pm$ 0.8 26.4 $\pm$ 0.8 27.3 $\pm$ 0.8 26.5 $\pm$ 0.7 25.7 $\pm$ 0.7
Oracle GP-UCB 48.3 $\pm$ 1.2 46.9 $\pm$ 1.1 48.4 $\pm$ 1.1 46.5 $\pm$ 1.0 45.6 $\pm$ 1.0
Table 2: Average total regret and $\pm 1$ standard error for the subspace experiment as $|P|$ increases.
Algorithm Subspaces, $|P|$
5 8 12 16
MAP GP-TS 87.2 $\pm$ 1.0 89.9 $\pm$ 1.1 89.1 $\pm$ 0.9 90.9 $\pm$ 1.2
HP-GP-TS 88.3 $\pm$ 0.9 88.8 $\pm$ 0.9 89.5 $\pm$ 0.9 90.8 $\pm$ 0.9
PE-GP-TS 177.1 $\pm$ 1.4 269.5 $\pm$ 1.9 344.7 $\pm$ 2.3 396.9 $\pm$ 2.5
PE-GP-UCB 389.0 $\pm$ 1.5 526.0 $\pm$ 1.8 622.4 $\pm$ 2.3 688.0 $\pm$ 2.7
Oracle GP-TS 86.0 $\pm$ 1.0 84.1 $\pm$ 0.9 84.6 $\pm$ 1.0 84.8 $\pm$ 1.0
Oracle GP-UCB 217.3 $\pm$ 1.0 218.2 $\pm$ 1.0 218.6 $\pm$ 1.0 218.9 $\pm$ 0.9
Figure 12: Median cumulative regret on Intel temperature data (top), PeMS speed data (middle) and PNW precipication data (bottom). Errorbars correspond to first and last decile.
(a) Intel
(b) PeMS
(c) PNW Precipication
Figure 13: Quantiles of the final cumulative regret on the real-world data experiments. The median is shown with a black line. The whiskers correspond to the 5th and 95th percentile and the lower and upper edges of the boxes show the first and third quantile.